OpenStudy (anonymous):

Solve the following system of equations: 4x - 2y - z = -5 x - 3y + 2z = 3 3x + y - 2z = -5 (0, 1, 3) (0, -1, 3) (0, 1, -3) (0, -1, -3)

OpenStudy (anonymous):

Any idea how to start?

OpenStudy (anonymous):

no, this is like a pretest so i have no idea!!

OpenStudy (anonymous):

Alright first thing, pick a variable you want to eliminate (x, y or z)

OpenStudy (anonymous):

x

OpenStudy (anonymous):

Cool, so now lets give each equation a label, so 1. 4x-2y-x=-5 2. x-3y+2z = 3 3. 3x+y-2z = -5 This will make it easier for us when we start cancelling everything out and isolating variables. Ok so now lets take 1 and 2 and eliminate x from it, can you do that for me?

OpenStudy (anonymous):

im trying i dont think im doing it right....

OpenStudy (anonymous):

yeah.. nope i cant figure it out...

OpenStudy (anonymous):

Don't worry, just tell me what you did/ doing, I don't judge, I'm here to help you learn :), if you make mistakes that's a good sign for me because I know you're trying, so just tell me what you did?

OpenStudy (anonymous):

Ok, well let me ask you this, we have 1. 4x-2y-z=-5, and 2. x-3y+2z=3, we want to make the x the same magnitude (number) but want the opposite sign for it so we can cancel it out, with that said, what would you do?

OpenStudy (anonymous):

sorry got caught up in feeding my baby sister.

OpenStudy (anonymous):

so i would take the x from the right and subtract it from the left right? @iambatman @$w3G_Godd Help!?

OpenStudy (anonymous):

@Gabebro13

OpenStudy (anonymous):

@ilovehim121511 since this is a pretest, I recommend answering randomly. You are not supposed to get it right if you don't know. So yeah

OpenStudy (anonymous):

@Gabebro13 Pre tests help you learn, that's the point. So you should treat it as if it's a test.

OpenStudy (anonymous):

Can you show me what you mean @ilovehim121511

OpenStudy (anonymous):

but it counts towards my grade... thats why im trying to get help, im not understanding....

OpenStudy (anonymous):

@iambatman @ilovehim121511 What I mean is you should try to answer the questions right, just that if you sincerely don't know, then you don't really need to worry about it.

OpenStudy (anonymous):

1. 4x-2y-z=-5, and 2. x-3y+2z=3 so it would be 3x-2y-z=-5 and -3y+2z+3 ?

OpenStudy (anonymous):

And the truth is I don't know how to solve this. Math isn't really my thing. If you need help with biology, just tag me.